OpenStudy (anonymous):

Simplify. (4x^2 + 7x - 5) - (3x^2 - 2x - 4)

OpenStudy (anonymous):

you can take each term on its own just distribute the minus sigh 4x^2 +7x - 5 - 3x^2 +2x +4 take the x^2, the x, and the constants seperatley 4x^2 - 3x^2 = x^2 7x +2x = 9x 4-5 = -1 so x^2 + 9x -1

OpenStudy (anonymous):

Thanks! :)